OpenStudy (anonymous):

A circle has an area of 841π cm2. What is the diameter?

OpenStudy (anonymous):

The formula for the area of a circle is \[A = \pi r^2\] so we get \[841 \pi = \pi r^2\] pi cancels and you get \[841 = r^2\] so\[r = \sqrt{841}\] and \[d=2r=2\sqrt{841}\] Make sense?

OpenStudy (anonymous):

Thank you!